# Searching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass Candidate

### v1

@@ 4.1 Alternative Light Curve Extraction
Figure 4: Detrended Sector 61 light curves for TIC-107150013, as extracted by TGLC (top) and QLP (bottom). The microlensing candidate at 2987 days is visually consistent between the two light curves, with similar shape, duration, and amplitude.
4.2 Lens Modeling
@@
We fit both QLP and TGLC light curves using point-source point-lens (PSPL) and finite-source point-lens (FSPL) models with linear limb-darkening, as implemented by the pyLIMA lensing model package (Bachelet et al., 2017). For the FSPL fits we used the pyLIMA FSPLargemodel, which is appropriate for the finite source magnification of large stars. We adopted a linear limb-darkening coefficient of 0.68 based on TESS values estimated by Claret (2017) for relatively cool giant stars ($\log g\sim 2.5$, $T_{\mathrm{eff}}\sim 4100$ K). We fixed the blending fraction to zero given that: (a) the light curves produced by the TGLC and QLP extractions have already undergone de-blending of resolved stars (§4.1), and (b) the nominal source is very bright in comparison to unresolved background stars, dominating the flux. Parallax effects due to the motion of the satellite are negligible on these timescales for all realistic lens masses and distances, [^5] One can estimate the order of magnitude of this effect by taking TESS’s displacement during the event to be $\Delta d\sim 25,000$ km and estimating $\pi_{E}\approx\Delta d/R_{E}$. This is $\lesssim 0.15$ for $M\sim 0.1M_{\oplus}$ and $d_{l}\sim 0.5\,\text{kpc}$ and only decreases with increasing mass and lens distance. and thus were not included in the fit. Ultimately, we performed three-parameter PSPL fits parameterized by $t_{0}$, $u_{0}$, and $t_{E}$, and four-parameter FSPL fits with the additional parameter $\rho$.
To explore the parameter space, we used the multinest Bayesian inference tool (Feroz et al., 2009) with 500 live points. We removed all data more than 2 days from $t_{\mathrm{BLS}}$ to speed up the fit. The TGLC and QLP fits converged after $\lesssim 25,000$ total likelihood computations, giving the results shown in Table 2 and Figure 5. multinest also provides the Bayesian log-evidence ($\log{Z}$) for each model, which can be used for model selection. The PSPL fits had significantly smaller $\log{Z}$ values compared to the FSPL fits ($\log{Z_{\mathrm{FSPL}}}-\log{Z}_{\mathrm{PSPL}}=12.0\pm 0.3$ and $7.8\pm 0.3$ for TGLC and QLP, respectively), indicating that the FSPL fits are overwhelmingly preferred. The TGLC corner plots for the FSPL fit are shown in Figure 6.
@@
Table 2: Best-fit PSPL and FSPL model parameters based on the TGLC and QLP Sector 61 light curves of TIC-107150013. The central values are the median of the posteriors, with lower and upper uncertainties from the 16th and 84th percentiles. We also report the Bayesian log-evidence values for each model, computed with multinest, indicating that the FSPL fits are preferred.
@@
$t_{0}$ $\mathcal{U}(2986.6,2987.4)$ $2987.017_{-0.002}^{+0.002}$ $2987.015_{-0.003}^{+0.003}$ $2987.017_{-0.002}^{+0.002}$ $2987.015_{-0.003}^{+0.003}$
$u_{0}$ $\mathcal{U}(0,10)$ $1.929_{-0.009}^{+0.010}$ $1.979_{-0.012}^{+0.013}$ $4.30_{-0.05}^{+0.05}$ $4.48_{-0.07}^{+0.08}$
$t_{E}$ $\mathcal{U}(0.01,1)$ $0.088_{-0.001}^{+0.001}$ $0.088_{-0.002}^{+0.002}$ $0.074_{-0.002}^{+0.002}$ $0.074_{-0.002}^{+0.002}$
$\rho$ $\mathcal{U}(5\times 10^{-5},100)$ $4.55_{-0.07}^{+0.08}$ $4.73_{-0.09}^{+0.12}$
$\log{Z}$ $-554.2\pm 0.2$ $-596.5\pm 0.2$ $-542.1\pm 0.2$ $-588.6\pm 0.2$
@@
Figure 5: The TESS microlensing candidate associated with the source star TIC-107150013, as seen in TGLC (left) and QLP (right) sector 61 light curves. The red lines show the median PSPL fits, while the black lines show the median FSPL model fits. Based on comparing $\log{Z}$ values for each result, the FSPL fits are overwhelmingly preferred.
@@
Figure 6: Corner plot displaying the multinest posteriors from the FSPL fit to the TGLC light curve of TIC-107150013.
The peak amplitude in the QLP light curve appears to be $\sim 0.4\%$ shallower than in the TGLC light curve, likely due to the more accurate dilution estimates in TGLC light curves compared to those of QLP. Consequently, we find FSPL $u_{0}$ and $\rho$ values that differ at the $2\sigma$ and $1.5\sigma$ level, respectively. However, both fits detect significant and similar finite-source effects, with the TGLC fit finding $\rho=4.55_{-0.07}^{+0.08}$ and the QLP fit finding $\rho=4.73_{-0.09}^{+0.12}$. We take the discrepancy in these two results as an indicator of systematic uncertainty associated with our extraction methodology.
4.3 Lens Mass and Velocity Constraints
The appearance of finite-source features in the light curve allows us to directly estimate the Einstein angle of the lens, which in turn provides constraints on the lens mass and relative velocity. Given TIC-107150013’s distance of $d_{s}=3193\pm 153$ pc and radius of $R_{s}=12.91_{-1.51}^{+0.96}~{}R_{\odot}$, its angular size is $\theta_{s}=9.11_{-1.15}^{+0.81}\times 10^{-11}$ radians. Assuming $\rho=4.55_{-0.07}^{+0.08}$ from the TGLC fit, this corresponds to an Einstein angle of $\theta_{E}=2.0_{-0.3}^{+0.2}\times 10^{-11}$ radians, or $4.1_{-0.5}^{+0.4}~{}\mu\mathrm{as}$. Solving Eq. (1) for the mass of the lens at a given lens distance, we found that our lens candidate is consistent with a planetary-mass object (Figure 7, blue curve). Similarly, $v_{\text{rel}}$ can be estimated from Eq. (2), which for $t_{E}=0.074$ days yields values on the order of 100 km/s (Figure 7, red curve). In Fig. 7, the shaded regions correspond to the uncertainty associated with measured values of $R_{s}$ and $d_{s}$ (Table 1). The formal uncertainties in the fit parameters for $\rho$ induce negligible changes to the lens mass and relative velocity estimates. Adopting instead the QLP value $\rho=4.73$ results in a uniform $\approx 7\%$ reduction in the mass estimates.
Though the nature of the lens cannot be directly measured through microlensing alone, these estimates provide strong motivation that the short-duration event associated with TIC-107150013 was due to a planetary-mass object moving at a velocity consistent with typical Galactic velocities, and we therefore label this event the first FFP candidate detected by TESS.
@@ 4.3 Lens Mass and Velocity Constraints
Figure 7: Lens masses (blue) and relative velocities (red) at different lens distances based on the FSPL fits to the TGLC light curve. The shaded regions correspond to the 16th to 84th percentiles of the fit results, with uncertainties from $R_{s}$ and $d_{s}$ propagated, while the solid line corresponds to the median values. The dotted grey line corresponds to 13 $M_{\mathrm{Jup}}$, which we assume is the largest possible planetary mass. At all lens distances, our candidate is consistent with a planetary mass object moving at typical Galactic velocities.
4.4 Consistency with Expected Yield

### v2

@@ 4.1.1 TESS-Gaia Light Curves
The re-detrended light curves revealed the presence of an additional small transient feature at $t=2972$ days that was below the threshold for significant detection in the original QLP extraction. As discussed in §3.3.3, microlensing events are expected to be isolated. Hence, the presence of this feature suggests some form of false positive contamination. In §4.2, we first operate under the assumption that the primary event is due to microlensing to highlight some of its more unusual aspects, then in §4.3 discuss this interpretation in the context of false positive contamination.
@@
Figure 7: Detrended Sector 61 light curves for TIC-107150013, as extracted by TGLC (top) and QLP (bottom). The black points show the data binned in 1-hr bins. The signal at 2987 days is visually consistent between the two light curves, with similar shape, duration, and amplitude.
4.1.2 SPOC Light Curves
We also obtained light curves for TIC-107150013 that were processed by the Science Processing Operations Center (SPOC; Jenkins et al., 2016) at the NASA Ames Research Center. SPOC produces light curves at shorter cadences than the FFIs for a pre-selected list of $\sim 20,000$ targets each sector. TIC-107150013 was not selected for SPOC processing in Sectors 7, 34, and 61, but we obtained 20-sec cadence observations in Sectors 87 and 88 through a Director’s Discretionary Targets (DDT) Proposal (DDT079, PI: Kunimoto). While these observations are long after the Sector 61 event, the 20-sec cadence observations are valuable for characterizing any potential flaring nature of the host star, where flares are potential microlensing false positives.
In Figure 8, we show the 20-sec cadence SPOC Pre-search Data Conditioning Simple Aperture Photometry (PDCSAP; Smith et al., 2012; Stumpe et al., 2012, 2014) light curves for Sectors 87 and 88, with all data points with a nonzero quality flag removed. Differences in the appearance of long-term variability compared to the QLP light curves (Figure 5) can be attributed to differences between light curve extraction procedures and systematics corrections; however, there are no repeated signals resembling the event from Sector 61, nor any indications of flares.
Figure 8: 20-sec cadence SPOC PDCSAP light curve for TIC-107150013 from Sectors 87 and $8$. There are no repeated signals resembling the event from Sector 61, nor any indications of flares.
4.2 Microlensing Interpretation
In this section, we treat the signal as a microlensing event. We fit the signal to a microlensing template and use the results to check the consistency of this event in three ways: (1) consistency with the expectation for a light curve produced by a low-mass FFP, (2) consistency of the implied mass and velocity with expectations for nearby FFPs, and (3) consistency of the implied event rate with the expected abundance of FFPs.
4.2.1 Consistency with Light Curve Morphology
@@
We fit both QLP and TGLC light curves using point-source point-lens (PSPL) and finite-source point-lens (FSPL) models with linear limb-darkening, as implemented by the pyLIMA lensing model package (Bachelet et al., 2017). For the FSPL fits we used the pyLIMA FSPLargemodel, which is appropriate for the finite-source magnification of large stars. We adopted a linear limb-darkening coefficient of 0.68 based on TESS values estimated by Claret (2017) for relatively cool giant stars ($\log g\sim 2.5$, assumed given the star’s giant nature; $T_{\mathrm{eff}}\sim 4100$ K, based on the star’s $T_{\mathrm{eff}}$ provided in the TICv8.2). We also fit for the flux contributions from both the source and possible blended stars ($f_{\mathrm{source}}$ and $f_{\mathrm{blend}}$, respectively). Parallax effects due to the motion of the satellite are negligible on these timescales for all realistic lens masses and distances and thus were not included in the fit.
Ultimately, we performed five-parameter PSPL fits parameterized by $t_{0}$, $u_{0}$, $t_{E}$, $f_{\mathrm{source}}$, and $f_{\mathrm{blend}}$, and six-parameter FSPL fits with the additional parameter $\rho$. We required $f_{\mathrm{source}}\geq 0$ and assigned a Gaussian prior to $f_{\mathrm{blend}}\sim\mathcal{N}(0,\sigma^{2})$ where $\sigma$ is 10 times the photometric uncertainty ($\sigma=55~e^{-}s^{-1}$ and $74~e^{-}s^{-1}$ for TGLC and QLP, respectively). The prior is centered on 0 based on our expectation that the blending fraction should be low, given that (a) the light curves produced by the TGLC and QLP extractions have already undergone de-blending of resolved stars (§4.1), (b) TIC-107150013 is very bright in comparison to unresolved background stars, dominating the flux, and (c) TIC-107150013 is consistent with being a single star based on Gaia DR3 astrometric data (RUWE = 0.99). We adopted uniform priors for all other fit parameters, as shown in Table 2.
To explore the parameter space, we used the emcee ensemble sampler (Foreman-Mackey et al., 2013; Goodman & Weare, 2010) with 50 walkers. Following Goodman & Weare (2010), we compute the integrated autocorrelation time for each fit parameter in order to assess convergence. We consider the chains converged if the number of steps is at least 50 times longer than the average autocorrelation time, and throw out burn-in steps equal to twice the maximum autocorrelation time. The TGLC light curve fits converged after 2700 and 3400 steps for PSPL and FSPL models, respectively, while the QLP light curve fits converged after 2800 steps and 3900 steps.
Our fit results are shown in Table 2 and Figure 9. We include the BIC values for each model in Table 2 and use this as our basis for model comparison. We chose BIC over alternative metrics such as the Akaike Information Criterion (AIC) because BIC is well-suited to comparing the explanatory capabilities of competing models (Shmueli, 2010). BIC imposes a heavier penalty than AIC for model complexity, which helps guard against overfitting. Despite the increased number of fit parameters required by the FSPL model, we find that PSPL fits had significantly larger BIC values ($\mathrm{BIC}_{\mathrm{PSPL}}-\mathrm{BIC}_{\mathrm{FSPL}}=$ $21.0$ and $12.0$ for TGLC and QLP, respectively), indicating that the FSPL fits are overwhelmingly preferred ($\Delta\mathrm{BIC}>10$; Kass & Raftery, 1995). The TGLC corner plots for the FSPL fit are shown in Figure 10. Both fits give blended flux contributions consistent with 0, and recover statistically significant finite source effects, with $\rho=4.57_{-0.15}^{+0.15}$ from TGLC and $\rho=4.75_{-0.23}^{+0.23}$ from QLP.
@@
Table 2: Best-fit PSPL and FSPL model parameters based on the TGLC and QLP Sector 61 light curves of TIC-107150013, including $\theta_{E}$ derived from the FSPL fit results and stellar angular size. The central values are the median of the posteriors, with lower and upper uncertainties from the 16th and 84th percentiles. We also report the BIC values for each model, indicating that the FSPL fits are preferred.
@@
$t_{0}$ (BJD - 2457000) $\mathcal{U}(2986.6,2987.4)$ $2987.017_{-0.002}^{+0.002}$ $2987.015_{-0.003}^{+0.003}$ $2987.017_{-0.002}^{+0.002}$ $2987.015_{-0.003}^{+0.003}$
$u_{0}$ $\mathcal{U}(0,10)$ $1.93_{-0.04}^{+0.04}$ $1.98_{-0.06}^{+0.06}$ $4.33_{-0.15}^{+0.15}$ $4.51_{-0.22}^{+0.21}$
$t_{E}$ (days) $\mathcal{U}(0.01,1)$ $8_{-0.002}^{+0.002} 0.0880.0$$8_{-0.003}^{+0.003}$ $074_{-0.002}^{+0.002} 0.0$$074_{-0.003}^{+0.003}$
$\rho$ $\mathcal{U}(5\times 10^{-5},10)$ - - $4.57_{-0.15}^{+0.15}$ $4.75_{-0.23}^{+0.23}$
$f_{\mathrm{source}}$ ($e^{-}s^{-1}$) $\geq 0$ $872_{-54}^{+54}$ $867_{-72}^{+73}$ $875_{-52}^{+52}$ $875_{-74}^{+72}$
[…]
$\theta_{E}$ ($\mu\mathrm{as}$) - - - $4.07_{-0.48}^{+0.43}$ $3.91_{-0.48}^{+0.45}$
BIC $56253.9$ $62318.7$ $56232.9$ $62306.7$
@@
Figure 9: The TESS event associated with the source star TIC-107150013, as seen in TGLC (left) and QLP (right) sector 61 light curves. The red lines show the median PSPL fits, while the black lines show the median FSPL model fits. Based on comparing BIC values for each result, the FSPL fits are overwhelmingly preferred. The residuals show the light curve after removal of the FSPL model.
@@
Figure 10: Corner plot displaying the emcee posteriors from the FSPL fit to the TGLC light curve of TIC-107150013.
There remains a small feature in the residuals around $t=2986.85$ days that is particularly visible in the TGLC light curve for the best-fit FSPL model. This may be due to systematic noise or a genuine brightening of the source. Beyond an FSPL model alone, we attempt to explain the event as a double flare in §4.3 and find that the FSPL interpretation is still strongly preferred. Hence, from these analyses, we conclude that the primary brightening event associated with this light curve is consistent with the expectations of a low-mass FFP microlensing event, making it impossible to eliminate as a false positive based on the local light curve morphology alone.
4.2.2 Consistency with Physical Parameters
Though the reconstruction of finite source features is challenging owing to degeneracies with other fit parameters (Johnson et al., 2022), the appearance of putative finite-source features in the light curve allows a direct estimate the Einstein angle that would be associated with the best-fit parameters, which in turn provides an estimate of the lens mass and relative velocity. We compare these to typical planetary masses and the local Galactic velocity distribution (Gaia Collaboration et al., 2018b), as false positives would likely produce physical parameter estimates inconsistent with the expectations for a population of FFPs. Given TIC-107150013’s distance of $d_{s}=3193\pm 153$ pc and physical radius $R_{s}=12.91_{-1.51}^{+0.96}~R_{\odot}$, we estimate its angular size as $\theta_{s}=R_{s}/d_{s}~\mathrm{rad}=18.8_{-2.1}^{+1.9}~\mu\mathrm{as}$. Assuming $\rho=4.57_{-0.15}^{+0.15}$ from the TGLC fit, this corresponds to an Einstein angle of $\theta_{E}=4.07_{-0.48}^{+0.43}~\mu\mathrm{as}$.
In the following, we again emphasize that the calculations are carried out under the working assumption of a detected event for the sake of clarity. For an event with measured $t_{E}$ and $\theta_{E}$, there remains an inherent degeneracy between the physical parameters $M,v_{T}$, and $d_{L}$. As such, in Fig. 11, we plot both the mass (blue) and transverse velocity (red) as a function of lens distance. Without additional information about the system, e.g. a measurement of $d_{L}$ from parallax, this degeneracy cannot be fully broken to make individual estimates of each parameter. However, one can take a Bayesian approach and build a posterior for the lens mass and distance by adopting a prior over the distribution of FFPs, accounting for the likelihood of recovering the observed values of $\theta_{E}$ and $t_{E}$. We take the reasonable prior on $d_{L}$ that the FFP abundance tracks the stellar abundance along the line of sight to TIC-107150013 using the exponential disk profile from Binney & Tremaine (2008) and assume a conservative log-uniform prior on $M$ over $10^{-3},10^{3}]~M_{\oplus}$. Assuming the lens and source velocities are independently drawn from a 3D Gaussian with a mean of 0 km/s and standard deviation $\sigma=55~\text{km/s}$, we calculate the $v_{T}$ by taking the vector difference and projecting onto the transverse plane. This results in a Rayleigh distribution for the transverse velocity
$P(v_{T})=\frac{v_{T}}{\sigma_{r}^{2}}\exp\Big(-\frac{v_{T}^{2}}{2\sigma_{r}^{2}}\Big),$
where $\sigma_{r}=\sqrt{\frac{3}{2}}\sigma$ and $v_{T}\geq 0$.
We use a Gaussian likelihood on $\theta_{E}$ and $t_{E}$ with standard deviation given by the mean of their upper and lower uncertainties. Performing the marginalization over this likelihood yields $M=0.53^{+0.40}_{-0.26}~M_{\oplus},d_{L}=675^{+359}_{-287}~\text{pc},v_{T}=63.26^{+31.79}_{-26.46}~\text{km/s}$ and is not strongly dependent on the assumed priors. For example, adopting a significantly different prior on the mass corresponding to the normalized predicted event rate (cf. Fig. 1) originating from the best-fit FFP mass function of Sumi et al. (2023) yields $M=0.33^{+0.46}_{-0.15}~M_{\oplus},d_{L}=490^{+425}_{-219}~\text{pc},v_{T}=43.80^{+40.93}_{-17.91}~\text{km/s}$.
It is also worth noting that before measuring $\theta_{E}$ and $t_{E}$, the predicted event rate alone indicates that TESS’s sensitivity to FFPs would peak at masses around $10^{-2},10^{-1}]~M_{\oplus}$. The recovered mass for the event in question is consistent with this expectation. However, this implicitly assumes that the mass function presented in Sumi et al. (2023) remains a good model of the FFP abundance for $M<M_{\oplus}$, a mass range that is currently unconstrained by observations. Thus the estimates associated with such a calculation are subject to considerable uncertainty, further complicating the interpretation of this particular event. Taken as a whole, this discussion demonstrates that under the assumption of a range of reasonable priors, the event is consistent with that of a low-mass planet traveling at a reasonable velocity relative to the local stellar distribution (Gaia Collaboration et al., 2018b), making it even more challenging to flag as a false positive.
@@ 4.2.2 Consistency with Physical Parameters
Figure 11: Hypothetical lens masses (blue) and relative velocities (red) at different lens distances based on the FSPL fits to the TGLC light curve. The shaded regions correspond to the 16th to 84th percentiles of the fit results for $\theta_{E}$ and $t_{E}$, with uncertainties from $R_{s}$ and $d_{s}$ propagated, while the solid line corresponds to the median values. Note that some regions are more likely than others, as discussed in 4.2.2. The dotted grey line corresponds to 13 $M_{\mathrm{Jup}}$, which we assume is the largest possible planetary mass.
4.2.3 Consistency with Expected Yield